A solid sphere of radius R has a nonuniform charge distribution p=Ar², where A is a constant. Determine the total charge, Q, within the volume of the sphere.

Please explain and show work

Thank you
Answer #1


volume element is dV = (4pi*r^2)*dr

then charge in volume element is dQ = rho*dV = rho*4*pi*r^2*dr

given that rho = A*r^2
then

dQ = 4pi*r^4*A*dr


total charge is Q = integral (dQ) = integral (A*4*pi*r^4*dr) = A*4*pi*integral(r^4*dr) = A*4*pi*(r^5/5) = (4/5)*A*pi*r^5 is the final answer